40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 13 12:21:09 2018, gautam, Update, ALS, PDFR laser checkout 
    Reply  Wed Jun 13 22:07:31 2018, gautam, Update, ALS, BeatMouth PDFR measurement IMG_7056.JPGBeatMouthPDFR.pdfBeatMouth_PDFRdata.tgz
       Reply  Fri Jun 15 14:22:05 2018, gautam, Update, ALS, BeatMouth PDFR measurement BeatMouthPDFR.pdfBeatMouth_PDFRdata.tgz
Message ID: 13957     Entry time: Wed Jun 13 22:07:31 2018     In reply to: 13955     Reply to this: 13973
Author: gautam 
Type: Update 
Category: ALS 
Subject: BeatMouth PDFR measurement 

Summary:

Neither of the Menlo FPD310 fiber coupled PDs in the beat mouth have an optoelectronic response (V/W) as advertised. This possibly indicates a damaged RF amplification stage inside the PD.

Motivation:

I have never been able to make the numbers work out for the amount of DC light I put on these PDs, and how much RF beat power I get out. Today, I decided to measure the PD response directly.

Details:

In the end, I decided that slightly modifying the Jenner laser setup was the way to go, instead of futzing around with the PDFR laser. These PDs have a switchable gain setting - for this measurement, both were set to the lower gain such that the expected optoelectronic response is 409 V/W.

[Attachment #1] - Sketch of the experimental setup. 

[Attachment #2] - Measured TF responses, the RF modulation was -20dBm for all curves. I varied the diode laser DC current a little to ensure I recovered identical transfer functions. Assumptions used in making these plots:

  1. NF1611 and FPD310 have equal amounts of power incident on them.
  2. The NF1611 transimpedance is 700V/A.

[Attachment #3] - Tarball of data + script used to make Attachment #2.

Conclusions:

  • The FPD310 does not have a DC monitor port. 
    • So the dominant uncertainty in these plots is that I don't know how much power was incident on the PD under test.
    • The NF1611 DC power level could be measured though, and seemed to scale with DC pump current linearly (I had only 3 datapoints though so this doesn't mean much).
  • Neither PD has transimpedance gain as per the specs.
    • The X PD shows levels ~x10 lower than expected.
    • The Y PD shows levels ~x3 lower than expected.
  • I will repeat the measurement tomorrow by eliminating some un-necessary patch fiber cables, and also calibrating out the cable delays.
    • The setup shown in Attachment #1 was used because I didn't want to open up the BeatMouth.
    • But I can pipe the port of the BS not going to the FPD310 directly to the collimator, and that should reduce the systematic uncertainty w.r.t. power distribution between FPD310 and NF1611.
Attachment 1: IMG_7056.JPG  1.395 MB  Uploaded Wed Jun 13 23:20:05 2018  | Hide | Hide all
IMG_7056.JPG
Attachment 2: BeatMouthPDFR.pdf  145 kB  Uploaded Wed Jun 13 23:24:21 2018  | Hide | Hide all
BeatMouthPDFR.pdf
Attachment 3: BeatMouth_PDFRdata.tgz  57 kB  Uploaded Wed Jun 13 23:24:28 2018
ELOG V3.1.3-